# mDP3: A Training-free Approach for List-wise Frame Selection in Video-LLMs

### v1

@@ 4.2.4 Ablation Study
+ mDP3 53.3
Table 3: Accuracies (%) of various frame extraction methods and ablation variants of mDP3 on Video-MME (without subtitles), based on VILA. Qry: whether considering query relevance; Div: whether considering list-wise diversity; Seq: whether considering list-wise sequentiality; HDS: whether employ high-dimensional similarity in RKHS; TFMA: whether tuning-free and model-agnostic for Video-LLMs.
@@
We conducted an ablation study of mDP3, with the results shown in Tab. 3 and Fig. 3. The study considered variations such as directly using SigLIP to select the top-$k$ similar frames based on the query, using standard DPP with CMGK while disregarding sequentiality, and applying mDP3 with similarity matrix modifications such as MGK and cosine similarity. We also compared mDP3 with traditional rule-based shot boundary detection methods, such as RGB Histogram [45], Edges Change Ratio [40], and Optical Flow [57], which select frames based on significant transitions between adjacent frames. These methods do not account for the current query or the list-wise relationship among selected frames. The cluster-based method Katna [17] only considers pair-wise diversity, overlooking query relevance, list-wise diversity, and sequentiality. Frame-Voyager [66] is a recent end-to-end frame selection method specifically designed for VILA. The ✓✗ symbol indicates that some factors, while not explicitly addressed, may be implicitly captured. Training on VILA’s intermediate hidden states with a list-based loss enables Frame-Voyager to potentially capture list-wise information, while position embeddings may address sequentiality.
@@ 5 Conclusion
This paper emphasizes that good frame selection should consider query relevance, list-wise diversity, and sequentiality. We propose mDP3, a training-free and model-agnostic method that fully addresses these principles for the first time. Experiments on three long video benchmarks show that integrating mDP3 with existing Video-LLMs significantly improves performance, demonstrating its effectiveness and robustness. mDP3 also provides a $(1-1/e)$-approximation for the NP-hard list-wise frame selection and selection size allocation problem in sequential video segments. This method has a practical time complexity of $\mathcal{O}(nk^{4})$, reducible to $\mathcal{O}(nk^{2})$ through lazy and parallel updates. Future work could use training-free mDP3 to refine the quality of training data, improving data efficiency.
@@ A.1 Preliminaries
A.2 Proof of $(1-1/e)$-approximation
Submodular maximization is NP-hard in general. Therefore, most research in this field focuses on approximation algorithms with polynomial-time complexity. While the submodular function is monotone, i.e., for every $X,Y\subseteq\Omega$, we have $f(X)\leq f(Y)$. The problem of maximizing a monotone submodular function subject to a cardinality constraint admits a $(1-1/e)$-approximation greedy algorithm (as introduced in Sec. 3.2[41].
In this section, we provide a concise proof of the $1-1/e$ approximation ratio for the greedy algorithm.
@@ A.2 Proof of $(1-1/e)$-approximation
$j=\mathop{\mathrm{argmax}}\limits_{i\in\Omega\setminus S}\quad f(S\cup\{i\})-f(S)\,.$ (25)
@@
This greedy algorithm provides a solution $\hat{S}$, which guarantees a $(1-1/e)$ approximation, where the optimal solution is denoted as $S^{*}$:
@@
$\exists x\in S^{*},\quad f_{S}(x)\geq\frac{1}{k}f_{S}(S^{*})\,.$ (28)
@@
$f(S_{t+1})-f(S_{t})\geq f(S_{t}\cup\{x\})-f(S_{t}),\quad\forall x\in\Omega\setminus S_{t}\,.$ (30)
@@ A.3 Proof of Time Complexity
The time complexity of Algo. 1 is closely $\mathcal{O}(nk^{3})$, where $n$ represents the number of candidate frames and $k$ denotes the selection size.
@@
As indicated by the pseudocode in Algo. 1, there are three loops, with the iteration sizes specified as follows:
@@
$\left\{\begin{array}[]{ll}\mathcal{O}(nm^{2}),&\text{if }k<m^{2/3}\\ \mathcal{O}(nk^{3}),&\text{if }m^{2/3}\leq k<m\\ \mathcal{O}(nm^{2}k)<\mathcal{O}(nk^{3}),&\text{if }k\geq m\\ \end{array}\right.$ (38)
In practice, the segment size $m$ is typically small, so in most practical cases, the time complexity of Algo. 1 is upper-bounded by $\mathcal{O}(nk^{3})$. ∎
@@
Additionally, the iteration in line 6 of Algo. 1 is independent, enabling parallel updates. This results in a time complexity of $\mathcal{O}(nk^{2})$, making its efficiency comparable to that of the standard DPP. Moreover, without lazy strategies and parallel updates, the time complexity is closely $\mathcal{O}(nk^{4})$ in most practical cases, and the proof is similar.
Appendix B Experiments Details
@@ B.2 Dataset Details
B.3 Implementation Details of Multi-kernel
We denote $\alpha_{u}=(h_{u}\sigma_{u})^{2}$ as a single hyperparameter. Following the multi-kernel maximum mean discrepancy (MK-MMD) framework [34], the optimal $\beta_{u}$ can be optimized using a quadratic program (QP). However, optimizing $\beta_{u}$ is orthogonal of this work. In the recent official implementation of Long et al. [34], average weights $\beta_{u}=1/U$ were employed, yielding good performance. Hence, we adopt average weights for $\beta_{u}$ and concentrate on configuring $\alpha_{u}$. Consistent with Long et al. [34], Sun and Li [48], for both $g,k\in\mathcal{K}$, we set $\alpha_{u}$ to $2^{i}$ where $i\in\{-3,-2,0,1,2\}$ and use an averaged ensemble of multiple Gaussian kernels. Additionally, the difference between $g$ and $k$ is modulated by $\lambda$.
B.4 Additional Parameters Analysis
@@
mDP3 is a training-free, model-agnostic method that leverages pretrained VLMs. In our experiments, we evaluate mDP3 using three primary baselines: VILA-V1.5-8B [28], MiniCPM-V2.6-7B [63], and LLaVA-OneVision-7B [21]. All three baselines integrate the vision encoder from SigLIP [67], so mDP3 only needs to introduce the additional parameter from the text encoder in SigLIP. The parameter scales for each baseline are reported in Tab. 4. The results indicate that the additional parameters from the text encoder amount to no more than 6% of the original MLLM scale, which is negligible. More importantly, these parameters are pretrained in VLMs and do not require tuning with the specific MLLMs.
B.5 Latency Comparison
Figure 4: Latency of various processes during inference using MiniCPM-V2.6 on Video-MME (without subtitles).
Our method, mDP3, is training-free, which eliminates any additional latency during the training phase. Accordingly, we report the average latency of various processes during inference with MiniCPM-V2.6 on Video-MME (without subtitles), as illustrated in Fig. 4. To ensure fairness, parallel optimization for frame selection is not employed. The baseline refers to MiniCPM-V2.6 processing all 128 candidate frames without applying frame selection. By selecting only 8 essential frames for input into the MLLMs, mDP3 achieves a significant speedup during the inference stage, offering nearly a 10-fold improvement. Although this is smaller than the theoretical $\mathcal{O}(n^{2})$ speedup of $(128/8)^{2}=256$ times, we attribute the discrepancy to the numerous engineering optimizations involved in LLM inference. Nonetheless, a 10-fold speedup represents a significant improvement. Furthermore, the yellow segment in Fig. 4, which represents the frame selection stage of mDP3, accounts for 32.8% of the total process. This indicates that, even when compared to the ablation of the mDP3 frame selection, the additional latency introduced by the method remains minimal, underscoring its efficiency during inference. Moreover, the latency of the image processing stage in VLMs is notable, suggesting that optimizing this stage could lead to significant speedups in future work.
B.6 Case Study
@@
We sample six representative cases from Video-MME [13], illustrated in Figs. 5, 6, 7, 8, 9 and 10, to compare different frame selection methods, including uniform sampling (marked by [Uncaptioned image] in the top right corner of frame) and top-$k$ query-frame matching with SigLIP [67] (marked by [Uncaptioned image]). The selection by mDP3 is marked by [Uncaptioned image]. Besides, the ground-truth of option is colored by green.
@@
    Failure of Counting Across Frames (Figs. 6 and 7): As shown in Figs. 6 and 7, when posed with questions such as, “What is the total number of bird species visible in the video?” or “How many different kinds of animal faces appear in this video?”, these counting questions differ from the example in Fig. 5. Unlike the case in Fig. 5, the counted items are distributed across different frames and do not appear in a single frame. To answer accurately, it is necessary to select frames that include various relevant counted items. However, the baseline method of uniform sampling fails to identify specific counted items, while the query-frame matching approach struggles to avoid duplication, often focusing repeatedly on frames with the same item. In contrast, the frame selection by mDP3 demonstrates greater diversity, allowing for the inclusion of different frames with various items, thereby aiding MLLMs in accurately counting across frames.

### v2

@@ 4.2.4 Ablation Study
[…]
Table 3: Results (%) of various frame selection methods and ablation variants of mDP3 on Video-MME (wosubswo\ subs), based on VILA. Indicators: query relevance (Qry), list-wise diversity (Div), list-wise sequentiality (Seq), high-dimensional RKHS similarity (HDS), and a tuning-free, model-agnostic design (TFMA).
@@
We conducted an ablation study of mDP3, with the results shown in Tab. 3 and Fig. 3. The study considered variations such as directly using SigLIP to select the top-$k$ similar frames based on the query, using standard DPP with CMGK while disregarding sequentiality, and applying mDP3 with similarity matrix modifications such as MGK and cosine similarity. We also compared mDP3 with traditional rule-based shot boundary detection methods, such as RGB Histogram [46], Edges Change Ratio [41], and Optical Flow [60], which select frames based on significant transitions between adjacent frames. These methods do not account for the current query or the list-wise relationship among selected frames. The cluster-based method Katna [18] only considers pair-wise diversity, overlooking query relevance, list-wise diversity, and sequentiality. Frame-Voyager  [70] and SeViLA  [68] are recent end-to-end frame-selection methods designed for a specific model. The ✓✗ symbol indicates that some factors, while not explicitly addressed, may be implicitly captured. Training on VILA’s hidden states with a list-based loss enables Frame-Voyager to potentially capture list-wise information, while position embeddings may address sequentiality.
@@ 5 Conclusion
This paper emphasizes that good frame selection should consider query relevance, list-wise diversity, and sequentiality. We propose mDP3, a training-free and model-agnostic method that fully addresses these principles for the first time. Experiments on three long video benchmarks show that integrating mDP3 with existing Video-LLMs significantly improves performance, demonstrating its effectiveness and robustness. mDP3 also provides a (11/e)(1-1/e)-approximation for the NP-hard list-wise frame selection and selection size allocation problem in sequential video segments. This method has a practical time complexity of $\mathcal{O}(nk^{4})$, reducible to $\mathcal{O}(nk^{2})$ through lazy and parallel updates. Future work could leverage training-free mDP3 to perform test-time computation scaling up, generating self-distilled data for iterative Video-LMMs refinement.
Acknowledgements
This paper is supported by NSFC (62076121) and Major Program (JD) of Hubei Province (2023BAA024). The authors would like to thank Hao-Yuan He for his helpful feedback on drafts of the paper.
@@ A.1 Preliminaries
A.2 Proof of (11/e)(1-1/e)-approximation
Submodular maximization is NP-hard in general. Therefore, most research in this field focuses on approximation algorithms with polynomial-time complexity. While the submodular function is monotone, i.e., for every $X,Y\subseteq\Omega$, we have $f(X)\leq f(Y)$. The problem of maximizing a monotone submodular function subject to a cardinality constraint admits a (11/e)(1-1/e)-approximation greedy algorithm (as introduced in Sec. 3.2[42].
In this section, we provide a concise proof of the 11/e1-1/e approximation ratio for the greedy algorithm.
@@ A.2 Proof of (11/e)(1-1/e)-approximation
$j=\mathop{\mathrm{argmax}}\limits_{i\in\Omega\setminus S}\qquad f(S\cup\{i\})-f(S)\,.$ (25)
@@
This greedy algorithm provides a solution $\hat{S}$, which guarantees a (11/e)(1-1/e) approximation, where the optimal solution is denoted as $S^{*}$:
@@
$\exists x\in S^{*},\qquad f_{S}(x)\geq\frac{1}{k}f_{S}(S^{*})\,.$ (28)
@@
$f(S_{t+1})-f(S_{t})\geq f(S_{t}\cup\{x\})-f(S_{t}),\qquad\forall x\in\Omega\setminus S_{t}\,.$ (30)
@@ A.3 Proof of Time Complexity
The time complexity of Algorithm 1 is closely $\mathcal{O}(nk^{3})$, where $n$ represents the number of candidate frames and $k$ denotes the selection size.
@@
As indicated by the pseudocode in Algorithm 1, there are three loops, with the iteration sizes specified as follows:
@@
[…]
In practice, the segment size $m$ is typically small, so in most practical cases, the time complexity of Algorithm 1 is upper-bounded by $\mathcal{O}(nk^{3})$. ∎
@@
Additionally, the iteration in line 6 of Algorithm 1 is independent, enabling parallel updates. This results in a time complexity of $\mathcal{O}(nk^{2})$, making its efficiency comparable to that of the standard DPP. Moreover, without lazy strategies and parallel updates, the time complexity is closely $\mathcal{O}(nk^{4})$ in most practical cases, and the proof is similar.
Appendix B Experiments Details and More Discussions
@@ B.2 Dataset Details
B.3 Implementation Details
Visual Encoder:
The primary baselines, LLaVA-OneVision-7B [22] and MiniCPM-V2.6-7B [66], both use SigLIP [71] as the visual encoder, which we also adopt as the VLM in Eq. 2 to avoid including excessive additional parameters. Additionally, since the context length of the text encoder in SigLIP is 64, potentially insufficient for the entire question, we split the text sequence into multiple sequences of equal length, each no longer than 64 tokens. We then extract multiple text embeddings and aggregate them into a final text embedding using pooling.
Multi-kernel:
We denote $\alpha_{u}=(h_{u}\sigma_{u})^{2}$ as a single hyperparameter. Following the multi-kernel maximum mean discrepancy (MK-MMD) framework [35], the optimal $\beta_{u}$ can be optimized using a quadratic program (QP). However, optimizing $\beta_{u}$ is orthogonal of this work. In the recent official implementation of Long et al. [35], average weights $\beta_{u}=1/U$ were employed, yielding good performance. Hence, we adopt average weights for $\beta_{u}$ and concentrate on configuring $\alpha_{u}$. Consistent with Long et al. [35], Sun and Li [49], for both $g,k\in\mathcal{K}$, we set $\alpha_{u}$ to $2^{i}$ where $i\in\{-3,-2,0,1,2\}$ and use an averaged ensemble of multiple Gaussian kernels. Besides, the difference between $g$ and $k$ is modulated by $\lambda$.
Hyper-parameters:
We set the trade-off hyperparameter $\lambda=0.2$ and the segment size $m=32$ for all tasks and benchmarks in mDP3, determined through cross-validation on a subset of the LLaVA-OneVision-mid [22] training set.
B.4 Hyperparameter Sensitivity Analysis
Figure 4: Hyper-parameter sensitivity with trade-off $\lambda$.
Figure 5: Hyper-parameter sensitivity with segment size $m$.
To evaluate the robustness of mDP3, we perform a sensitivity analysis of the score-trade-off parameter $\lambda$ and the segment size $m$ using the VideoMME benchmark (without subtitles). The results are shown in Figs. 4 and 5.
Score-trade-off ($\boldsymbol{\lambda}$).
Figure 4 varies $\lambda$ logarithmically from 0.0125 to 12.8. Accuracy stays within a narrow range of 51–53 %, peaking at 53.3 % when $\lambda=0.2$; every value surpasses the baseline of 49.2 %.
Segment size ($\boldsymbol{m}$).
Figure 5 evaluates segment sizes ranging from 8 to 128 frames. All configurations exceed the baseline. The best accuracy, 53.3 %, is achieved at $m=32$; accuracy stays above 52 % for $m\leq 64$ and remains competitive (51.8 %) even at $m=128$. These findings indicate that mDP3 is insensitive to the choice of $m$.
Across all examined values of $\lambda$ and $m$, mDP3 consistently outperforms uniform sampling, confirming its robustness and practical ease of tuning.
B.5 Additional Parameters Analysis
@@
mDP3is a training-free, model-agnostic method that leverages pretrained VLMs. The primary baselines, VILA-V1.5-8B [29], MiniCPM-V2.6-7B [66], and LLaVA-OneVision-7B [22], all integrate the vision encoder from SigLIP [71], so mDP3 only needs to introduce the additional parameter from the text encoder in SigLIP. The parameter scales of them are reported in Tab. 4. The results indicate that the additional parameters from the text encoder amount to no more than 6% of the original MLLM scale, which is negligible. More importantly, these parameters are pretrained in VLMs and do not require tuning with the specific MLLMs.
B.6 Latency Comparison
Figure 6: Latency during inference with MiniCPM-V2.6 on Video-MME (without subtitles). Where mDP3-P refers o the mDP3 with parallel computation in dynamic programming (i.e. line 9 in Algorithm 1). For direct comparison and better visualiztion, we omit the latency of identical processes across all compared models, including image loading and processing in the visual processor and encoder.
Our method, mDP3, is training-free, which eliminates any additional latency during the training phase. Accordingly, we report the average latency of various processes during inference with MiniCPM-V2.6 on Video-MME (without subtitles), as illustrated in Fig. 6. The baseline refers to MiniCPM-V2.6 processing all 128 candidate frames without applying frame selection. Compared to the baseline, mDP3 selects only 8 essential frames as input to the MLLMs, achieving a significant speedup during inference. It reduces MLLM inference time by 11.47s (90% of the baseline) while requiring only an additional 1.94s (15% of the baseline) for frame selection. mDP3-P accelerates frame selection using parallel computation in dynamic programming (i.e. line 9 in Algorithm 1), requiring only an additional 1.30s for frame selection while reducing MLLM inference time by 11.47s (90% of the baseline). Furthermore, mDP3-P has a latency comparable to the basic DPP method, which accounts only for list-wise diversity. This accords to our theoretical analysis, as both mDP3-P and DPP share the same time complexity of $\mathcal{O}(nk^{2})$.
B.7 Additional Experiments about Ablation Study
We conduct detailed ablation experiments on three Video-LLMs (VILA-V1.5, MiniCPM-V2.6, LLaVA-OneVision) and three benchmarks (Video-MME, MLVU, LVBval), demonstrating the consistent superiority of mDP3 over alternative strategies.
mDP3 attains peak scores across all configurations: 53.3/58.6/50.8 for VILA-V1.5, 58.0/66.6/57.1 for MiniCPM-V2.6, and 59.6/69.8/59.0 for LLaVA-OneVision. The most significant improvements are observed on MLVU, with gains of +4.8 to +6.2 over the baselines, underscoring its effectiveness in complex video understanding tasks.
Baseline uniform sampling underperforms by 3.0 to 9.5 points across all models, highlighting the importance of frame selection. Furthermore, intermediate variants exhibit limited gains due to partial adherence to our proposed principles; neglecting query relevance, list-wise diversity, or sequentiality fundamentally restricts selection quality. Moreover, the ablation results further validate the design choices of mDP3. For example, although the cosine similarity variant outperforms uniform sampling by 2.1 to 4.8 points, it still lags behind mDP3 by the same margin. This difference underscores the advantage of our multi-kernel similarity in the reproducing kernel Hilbert space (RKHS), which better captures high-dimensional feature relationships.
Video-MME MLVU LVBval
VILA-V1.5 + uniform 47.5 46.3 47.1
+ SigLIP 50.6 53.9 46.4
+ mDP3 w. MGK 48.9 48.6 53.5
+ DPP w. CMGK 51.8 56.8 47.1
+ mDP3 w. cos. sim. 50.2 54.4 48.8
[…]
MiniCPM-V2.6 + uniform 52.6 55.4 51.2
+ SigLIP 56.3 60.3 51.4
+ mDP3 w. MGK 52.6 62.0 52.6
+ mDP3 w. cos. sim. 53.1 63.0 54.1
+ DPP w. CMGK 55.2 64.7 52.1
[…]
LLaVA-OV + uniform 53.6 59.3 54.2
+ SigLIP 57.0 62.7 51.6
+ mDP3 w. MGK 54.9 63.3 52.7
+ DPP w. CMGK 56.4 68.6 52.8
+ mDP3 w. cos. sim. 55.9 65.2 54.6
[…]
Table 5: Ablation study of mDP3 across three Video-LLMs (VILA-8B, MiniCPM-V2.6, LLaVA-OneVision) and benchmarks (Video-MME, MLVU, LVBval).
B.8 More Experiments with Various Selection Size
Figure 7: Detailed accuracy (%) on Video-MME (without subtitles) for VILA-V1.5 grouped by various durations, comparing uniform sampling with mDP3 selection from 128 candidate frames.
Figure 8: Detailed accuracy (%) on Video-MME (without subtitles) for LLaVA-OneVision grouped by various durations, comparing uniform sampling with mDP3 selection from 128 candidate frames.
We have reported the results for various selection sizes $k$ in Sec. 4.2.2 and Fig. 1. In this section, we provide additional results and more detailed discussions. In Figs. 8 and 7, we present the performance of VILA-V1.5 and LLaVA-OneVision on Video-MME (without subtitles) across various durations by varying the selection size $k$. The results show that, regardless of duration, VILA-V1.5’s performance improves with more input frames when $k\leq 16$, while LLaVA-OneVision’s performance improves with additional frames when $k\leq 32$. They employ the distinct context lengths and numbers of frames used during training. When the number of input frames exceeds the training stage, the input becomes out-of-distribution (OOD), leading to a performance drop. Specifically, when VILA-V1.5 receives more than 16 input frames, it experiences a catastrophic decline, whereas LLaVA-OneVision maintains some generalization ability when presented with more than 32 frames. Although some Video-LLMs are trained with more frames to enhance long-video understanding, this approach is costly and the number of input frames cannot be increased indefinitely. Therefore, designing an effective frame selection algorithm is crucial.
There should be a meaningful discussion about the question: “Are more input frames better for video understanding?” In Tab. 1, we select 8 frames following the state-of-the-art frame selection approach, Frame-Voyage, settings to ensure a fair comparison. However, 8 frames may not be optimal for all Video-LLMs, benchmarks, and durations. Our answer to the question “Are more input frames better for video understanding?” is No! We provide the following reasons:
  1. 1.
    Training Constraints: Video-LLMs are typically trained on a limited number of frames. When the number of input frames exceeds the training configuration, the model encounters OOD data, leading to degraded performance. Although training with more frames can mitigate this issue, it is computationally expensive and unsustainable in the long run (i.e., the number of input frames cannot be increased indefinitely.).
  2. 2.
    Inference Constraints: Edge-deployed LLMs have limited resources, and proprietary models charge based on token usage. Consequently, processing excessive frames during inference stage is both resource-intensive and expensive.
  3. 3.
    Diminishing Returns: The marginal gain from adding frames diminishes exponentially (as indicated by the theoretical bound $\mathcal{O}((1/k)^{k})$ in Eq. 35 and supported by experiments in Figs. 1, 8 and 7), while the Transformer’s computational cost increases quadratically ($\mathcal{O}(n^{2})$). Consequently, adding more frames is not cost-efficient.
  4. 4.
    Redundancy and Noise: Incorporating additional frames may introduce redundant or irrelevant information, which can dilute salient features and add noise.
  5. 5.
    Latency and Real-Time Constraints: In applications requiring real-time processing, increasing the number of frames can lead to higher latency, which may be unacceptable in time-sensitive scenarios (as shown in Fig. 6).
These considerations underscore the importance of developing an effective frame selection algorithm that balances performance improvements with computational efficiency and resource limitations. Instead of merely increasing the number of input frames, a well-designed selection strategy can extract the most informative frames, ensuring that the model focuses on quality over quantity and optimizes both accuracy and cost.
Additionally, the above discussion is supported across various durations, encompassing short, medium, and long video understanding scenarios.
B.9 Qualitative Analysis
@@
We sample six representative cases from Video-MME [13], illustrated in Figs. 9, 10, 11, 12, 13 and 14, to compare different frame selection methods, including uniform sampling (marked by [Uncaptioned image] in the top right corner of frame) and top-$k$ query-frame matching with SigLIP [71] (marked by [Uncaptioned image]). The selection by mDP3 is marked by [Uncaptioned image]. Besides, the ground-truth of option is colored by green.
@@
    Failure of Counting Across Frames (Figs. 10 and 11): As shown in Figs. 10 and 11, when posed with questions such as, “What is the total number of bird species visible in the video?” or “How many different kinds of animal faces appear in this video?”, these counting questions differ from the example in Fig. 9. Unlike the case in Fig. 9, the counted items are distributed across different frames and do not appear in a single frame. To answer accurately, it is necessary to select frames that include various relevant counted items. However, the baseline method of uniform sampling fails to identify specific counted items, while the query-frame matching approach struggles to avoid duplication, often focusing repeatedly on frames with the same item. In contrast, the frame selection by mDP3 demonstrates greater diversity, allowing for the inclusion of different frames with various items, thereby aiding MLLMs in accurately counting across frames.